OpenStudy (lgbasallote):

how to prove L{t^n} = \(\large \frac{n!}{s^{n+1}}\) using the definite for laplace transform?

OpenStudy (anonymous):

ok are you familiar with Gamma function? because it will use its definition?

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

gamma function?

OpenStudy (anonymous):

yes .

OpenStudy (lgbasallote):

i dont think i am...not sure though...im not really familiar with fancy terms

OpenStudy (anonymous):

ok let me solve this first.

OpenStudy (anonymous):

first check gamma function here . http://en.wikipedia.org/wiki/Gamma_function

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

\[\text{ \let } I_n=\int\limits_{0}^{+\infty}t^{n}e^{-st}dt\] let \[n \in \mathbb{N}^{*}\] using an integration by parts \[I_n=\int\limits_{0}^{+\infty}(-\frac{e^{-st}}{s}) \prime t^n dt =\left[ (-\frac{e^{-st}}{s}) \prime t^n \right]"=0"+\frac{n}{s}\int\limits_{0}^{+\infty}t^{n-1}e^{-st}dt\] then \[I_n=\frac{n}{s}I_{n-1}\] (**) now there is two methods to do this : 1)\[I_n=\frac{n}{s}I_{n-1}=\frac{n}{s}\frac{(n-1)}{s}I_{n-2}=\frac{n}{s}\frac{(n-1)}{s}\frac{(n-2)}{s}I_{n-3}\]........................... \[I_n=\frac{n}{s}\frac{(n-1)}{s}\frac{(n-2)}{s}.....................\frac{1}{s}I_0=\frac{n!}{s^n}I_0\] if you calculate I_0 you will find 1/s 2) using induction for n=0 \[I_0=\frac{0!}{s^{0+1}}\] so the "statement" is true for n=0 let n be a positive integer let's assume that \[I_n=\frac{n!}{s^{n+1}}\] let's prove that \[I_{n+1}=\frac{(n+1)!}{s^{n+1+1=n+2}}\] using(**) \[I_{n+1}=\frac{(n+1)}{s}I_n=\frac{(n+1)}{s}\frac{n!}{s^{n+1}}=\frac{(n+1)!}{s^{n+2}}\] so we proved that \[\forall n \in \mathbb{N} I_n=L(t^n)=\frac{n!}{s^{n+2}}\] **"I don't like the first one" !

OpenStudy (anonymous):

@Neemo i think in the denominator it should be \[\Large S^{n+1}\] not n+2 !

OpenStudy (anonymous):

yes ! you're right !I made a mistake in the last line (n+1) How do I correct it .?

OpenStudy (anonymous):

@sami-21

OpenStudy (anonymous):

i am not Quite good with the Latex codes . you have to write it again:( or take snap till where the answer is correct. then write the remainingg:P

OpenStudy (anonymous):

I hope that @lgbasallote will read your comment ! or this "I made a typo in the last line" the correction " \[I_n=L(t^n)=\frac{n!}{s^{n+1}}\]" not n+2 :D

OpenStudy (anonymous):

there was typo in response as well :P corrected below.

OpenStudy (lgbasallote):

i have no idea what these mean =_=

OpenStudy (anonymous):

all of these !!!!! :p :p :p :p ???

OpenStudy (anonymous):

then take it :P before understanding gamma function you will not be able to do this :) so remember for time being just like ! \[\Large (a+b)^2=a^2+b^2+2ab\] so is \[\Large L(t^n)=\frac{n!}{s^{n+1}}\] :)